92. Investing her bonus. Donna invested her $33,000 bonus
and received a total of $970 in interest after one year. If
part of the money returned 4% and the remainder 2.25%,
then how much did she invest at each rate?

Answer by RAY100(1637)

let x+y=33000,,,,x=33000-y
where x & y are amounts invested
970=.04x+.0225y
970=.04(33000-y)+.0225y=1320-.04y-.0225y
970=1320-.0175y
-350=-.0175y
y=20,000
x-33000-y=33000-20000=13000
check,,,13000(.04)+20000(.0225)=520+450=970 ok